Total Assets $7,082,500
Total Liabilities 1,700,000
Common Stock 1,250,000
Additional Paid in Capital 2,097,500
Donated Capital 90,000
Ret. Earnings 1/1/Year4 1,650,000
Net Sales 6,250,000
Cost of Sales 3,750,000
Selling & Adm Expenses 1,212,500
Interest Expense 122,500
gain on sale of LT Investments 130,000
Income Tax Expense 300,000
Loss on Disposition of Plant Assets 225,000
Loss due to Earthquake Damage 475,000

Pucket Corp. is in the process of preparing its financial statements for the year ended December 31, Year4. Before closing the books, it prepared the above Condensed Trial Balance Sheet.

Other financial data for the year ended December 31, Year 4:

- Sales returns and allowances equaled $215,000, and sales discounts taken were $95,000.

- Estimated federal income tax payments were $200,000 and accrued federal income taxes equaled $100,000. The total charged to income tax expense does not properly reflect current or deferred income ta expense or interperiod income tax allocation for income statement purposes. The enacted tax rate on all types of taxable income for the current and future years is 30%. The alternative minimum tax is less than the regular income tax.

- Interest expense includes 6% interest on 20 year bonds issued at their face amount of $1,500,000.

- A $90,000 excess of carrying amount over tax basis in depreciable assets arose from receipt of a contribution of equipment by a local government on December 31, Year 4. it is expected to be depreciated over 5 years beginning in Year 5. There were no temporary differences prior to Year 5.

- Officer's Life insurance expense (not tax deductible) is $70,000.

- The earthquake damage is considered unusual and infrequent, but the disposition of plant assets is considered infrequent but not unusual. Moreover, the disposition of plant assets was not a disposal of a component of an entity.

- The shares of common stock ($5 par) traded on a national exchange:

Outstanding at 1/1/Year 4 200,000
Issued on 3/30 Year 4 as 10% Stock Dividend 20,000
Issued Shares for $25 per share on 6/30/Year 4 30,000
Outstanding at 12/31/Year 4 250,000

- Puckett declared a $1.25 common stock dividend on December 28, Year 4.

Using this information enter the correct amounts for Pucket Corporation's income statement for the year ended December 31, Year 4.

Net Sales
Cost of Sales
Gross Profit
Selling & Administrative Expenses
Income from Operations
Other Revenues and Gains:
Gain on Sale of LT Investments
Other Expenses and Losses:
Interest Expense
Loss on Disposition of Plant Assets
Income from continuing operations before income tax
Income Tax Expense:
Current Tax Expense
Deferred Tax Expense
Income Before Extraordinary Item
Extraordinary item-loss from Earthquake(net of applicable taxes)
Net Income

(if you could show all calculations as well that would be awesome!)
Answer #1
Pucket Corporation's income statement for the year ended December 31, Year 4.
Particulars Working Note ABBREVATION Amount ( In $)
Net Sales 1 A 5940000
Cost of Sales B 3750000
Gross Profit C=(A-B) 2190000
Selling & Administrative Expenses D 1212500
Income from Operations E=(C-D) 977500
Other Revenues and Gains: F 0
Gain on Sale of LT Investments G 130000
Other Expenses and Losses: H 70000
Interest Expense 2 I 95500
Loss on Disposition of Plant Assets J 225000
Income from continuing operations before income tax K=(E+G-H-I-J) 717000
Income Tax Expense: L=(M+N) 228000
Current Tax Expense M 207000
Deferred Tax Expense 3 N 21000
Income Before Extraordinary Item O=(K-L) 489000
Extraordinary item-loss from Earthquake(net of applicable taxes) P 332500
Net Income Q=(O-P) 156500
WORKING NOTE - 1
Net Sales
As per information provided 6250000
Adjustments
Sales returns and allowances -215000
sales discounts -95000
Total 5940000
WORKING NOTE - 2
Interest Expense
As per information provided 122500
Interest On Bonds 90000
1500000*6%
Tax (90000*30%) 27000
Interest Net Of Taxes 63000
Other Interest Expenses 32500
(122500-90000)
Total Interest Net Of Tax 95500
WORKING NOTE - 3
Deferred Tax Expense
Officer's Life insurance expense 70000
Rate of Income Tax 30%
Deferred Tax Expense (70000*30%) 21000
Total 21000
EXPLANATION
Receipt Of $90000 From Local Government As A Part Of Contribution Of Equipment Woul Be Deducted From
Cost Of Total Assets As A Part Of Capital Receipt As Per Accounting For Government Grant Principles Moreover
It Is Stated That The Equipment Woul Be Expected To Be Depriciated Over 5 Years Begining From The Fifth Year.
Hence While Preparing The Above Income Statement The Effect Of The Same Has Not Been Considered.
It Has Been Assumed That Effect Of Officers Life Insurance Has Not Been Provided Hence For Preparing Net Income
Statement Effect Of The Same Has Been Given Under Other Expenses.